OpenStudy (fanduekisses):

Let f(x)=sqrt(4x-3) and g(x)=f(x)/x What is the slope of the graph of f at x=3?

OpenStudy (fanduekisses):

Does that mean I have to find the derivative?

OpenStudy (amistre64):

yes

OpenStudy (fanduekisses):

\[f(x)= \sqrt{4x-3}\] \[g(x)=f(x)/x\]

OpenStudy (fanduekisses):

so is it zero?

OpenStudy (amistre64):

what is f'(x) ?

OpenStudy (fanduekisses):

2x^-1/2

OpenStudy (amistre64):

not quite

OpenStudy (fanduekisses):

but do I have to plug in the 3

OpenStudy (fanduekisses):

first?

OpenStudy (amistre64):

sqrt(u) to u' / 2sqrt(u)

OpenStudy (amistre64):

u' = 4 soo f' = 2/sqrt(4x-3)

OpenStudy (fanduekisses):

wait, what? what is 'sqrt(u) to u' / 2sqrt(u)'

OpenStudy (amistre64):

you know what a chain rule is right?

OpenStudy (fanduekisses):

nope :( I don't think I'm there yet

OpenStudy (amistre64):

well then youll have to work it with th elimit rule then

OpenStudy (fanduekisses):

I only know about the slope of tangent line (kinda-ish) the quotient and the power rule

OpenStudy (amistre64):

if we have a composition of fuctions, one function inside of another, then we apply what is called a chain rule h(t) derives to h'(t(x)) * t'(x) * x' but by convention: x' = dx/dx = 1

OpenStudy (amistre64):

right now we have a sqrt function with a linear function side of it sqrt(u) , such that u = 4x-3 the derivative is therefore: [sqrt(u)]' * u'

OpenStudy (amistre64):

*inside of it

OpenStudy (fanduekisses):

so \[\sqrt{4x-3} * 4\] ?

OpenStudy (amistre64):

almost, the first part needs a derivative

OpenStudy (amistre64):

\[\frac d{dx} \sqrt{4x-3}\] \[\frac d{dx} (4x-3)^{1/2}\] well, if this was sqrt(x) we would be able to define it, so lets define the innards by 1 variable, let u = 4x-3 \[\frac d{dx} (u)^{1/2}=\frac{du}{dx}\frac{1}{2u^{1/2}}\] but what is du/dx? we solve that from our definition of u \[\frac d{dx}u = \frac d{dx}4x-3\]\[\frac {du}{dx} = \frac {dx}{dx}4\] therefore du/dx=4 \[\frac d{dx} (u)^{1/2}=\frac{4}{2u^{1/2}}\]

OpenStudy (fanduekisses):

So I get \[8x^{-1/2}\]

OpenStudy (amistre64):

\[\frac{4}{2u^{1/2}}=2u^{-1/2}=2(4x-3)^{-1/2}\]

OpenStudy (amistre64):

let x=3 \[2(4(3)-3)^{-1/2}\] \[2(9)^{-1/2}\] \[2(3)^{-1}=2/3\]

OpenStudy (fanduekisses):

Ohhh I get it now ^_^ thanks

OpenStudy (amistre64):

yw